Solve using square root property
2x^2-35=15
Thanks in advance!!

Answer by checkley71(8403)

2X^2-35=15
2X^2-35-15=0
2X^2-50=0
2(X^2-25)=0
2(X-5)(X+5)=0
X-5=0
X=5 ANSWER.
X+5=0
X=-5 ANSWER.